# DySLIM: Dynamics Stable Learning by Invariant Measure for Chaotic Systems

### v1

@@ Appendix A Relation to the Pushforward trick
We point out that the pushforward trick can be reformulated using our framework as a weak measure fitting loss. A finite-sample approximation of $\mathbb{E}_{{\bm{u}}_{0}\sim\mu^{*}}[||{\mathcal{S}}_{\theta}(\mathtt{sg}({\mathcal{S}}^{k-1}_{\theta}({\bm{u}}_{0})))-{\bm{u}}_{k}||^{2}]$ is an upper bound of the discrete Wasserstein-2 distance between $\mu^{\star}$ and the approximation of $\mu_{\theta}^{\star}$. Formally, we have that
@@
By relying on an estimate of the loss, we have that for a given set of initial conditions $\{{\bm{u}}^{(i)}\}_{i=1}^{N}\sim\mu^{*}$ , we have that
@@
which can lower bounded following
@@
where $\pi$ is a permutation operator. Given that we are in the discrete setting both the Monge and Kantorovich problems are equivalent (Brezis, 2018), we have that
@@
where $\Pi$ is the set of all valid discrete transport maps, i.e., matrices that satisfy $T_{i,j}\geq 0$, $\sum_{j}T_{i,j}=\sum_{i}T_{i,j}=1$, where $C$ is the quadratic cost function $C_{i,j}=\|{\mathcal{S}}^{k}({\bm{u}}^{(i)})-{\mathcal{S}}_{\theta}^{k}({\bm{u}}^{(j)})\|^{2}$, and $\widehat{W}_{2}$ is a discrete estimate of the Wasserstein-2 metric.
@@ Appendix B Maximum Mean Discrepancy
In this section, we provide additional information and context about the Maximum Mean Discrepancy (MMD). The MMD is an instance of an integral probability metric (IPM) (Müller, 1997), which is a useful construction that allows us to measure distance between distributions. For any two distributions $\mu$ and $\nu,$ IPMs are defined with a function class ${\mathcal{G}}$ as:
@@
As described in Section 5, we use a rational quadratic kernel. While other works, such as Li et al. (2015) and Dziugaite et al. (2015), that use MMD for distribution matching also explored the squared exponential kernel, $\kappa_{\sigma}({\bm{u}},{\bm{v}})=\exp(\frac{-1}{2\sigma}||{\bm{u}}-{\bm{v}}||_{2}^{2})$, they found that careful tuning of the bandwidth parameter was required. In contrast, other than the highest dimension Kolmogorov flow experiments, we found that the mixture of bandwidths used in our rational kernel was comparatively robust and did not require a comprehensive hyperparameter search. We therefore rely on this kernel and do not explore the more sensitive squared exponential kernel.
Additionally, we note that we also explored using the Sinkhorn Divergence as the measure distance in Equation 13. However, especially in higher dimension experiments, we found this divergence to be less informative in guiding training, likely owing to its less favorable estimation properties compared to the MMD, particularly in the high-dimensional regime.
@@ C.1 Cosine Similarity
Letting $\{{\bm{u}}^{(i)}_{t_{k}}\}_{i=1}^{n}$ and $\{\tilde{{\bm{u}}}^{(i)}_{t_{k}}\}_{i=1}^{n}$ be the ground truth and predicted states at time $t_{k}$, for $k=1,...,N$, across test set trajectories, respectively, the cosine similarity at each time step is defined as:
@@ C.2 Sinkhorn Divergence
One of the popular metrics used to measure distance between distributions are the Optimal Transport (OT) based metrics, such as the Sinkhorn divergence, which we describe below. The field of OT is concerned with transforming (or transporting) one distribution into another, i.e., finding a map between them, in an optimal manner with respect to a pre-defined cost. The cost of the minimal (or optimal) transformation, often called the cost of the OT map, can then be used to define distances between distributions that ‘lifts’ the underlying metric $\mathrm{d}$ defined on ${\mathcal{U}}$ to one over the space of probability measures defined on ${\mathcal{P}}({\mathcal{U}})$ (Santambrogio, 2015).
@@
In Section 5, the SD was used to compare empirical version of the ground truth and predicted distributions of trajectories. We use the Optimal Transport Tools library (Cuturi et al., 2022) with its default hyperparameters to perform this computation.
@@ C.3 Radially Averaged Energy Spectrum
The energy spectrum is one of the main metrics to quantitatively assess the quality of the resulting snapshots (Wan et al., 2023a). In a nutshell, the energy spectrum measures the energy in each Fourier mode and thereby providing insights into the similarity between the generated and reference samples.
The energy spectrum is defined [^7] This definition is applied to each sample and averaged to obtain the metric (same for MELR below). as
@@
where ${\bm{u}}$ is a snapshot system state, and $k$ is the magnitude of the wave-number (wave-vector in 2D) $\underline{k}$, and $x_{i,j}$ is the underlying (possibly 2D) spatial grid. To assess the overall consistency of the spectrum between the generated and reference samples using a single scalar measure, we consider the mean energy log ratio (MELR):
@@ C.4 Covariance RMSE (covRMSE)
$\text{covRMSE}=\frac{1}{N}\sum_{n=1}^{N}\frac{\|\text{Cov}_{\text{pred}}-\text{Cov}_{\text{ref}}\|}{\|\text{Cov}_{\text{ref}}\|},$ (25)
@@ C.5 Time Correlation Metric (TCM)
can be interpreted as the time for the signal to forget its past. We compute the average pixel-wise $\tau$ for ground truth as well as prediction rollouts. To obtain a proper metric, we take the absolute difference from ground truth.
@@ Appendix D Regularization
To enhance the reproducibility of this paper, we provide explicit formulas for the objective functions. We reproduce Equation 13 from Section 5
@@ Appendix E Experimental Setup
We provide information about each dynamical system benchmarked and theird corresponding experimental setup.
@@ E.1 Lorenz 63
The Lorenz 63 model (Lorenz, 1963) is defined on a 3-dimensional state space by the following non-linear ordinary differential equation (ODE) $\dot{{\bm{u}}}=f({\bm{u}})$:
@@
The Lorenz 63 system is typically associated to parameter values of $\sigma=10,\rho=28,$ and $\beta=8/3$ and is known to be chaotic with an attractor that supports an ergodic measure (Luzzatto et al., 2005; Tucker, 2002).
@@
We parameterize ${\mathcal{S}}_{\theta}$ as a one step finite difference model: $\tilde{{\bm{u}}}_{k+1}={\mathcal{S}}_{\theta}({\bm{u}}_{k})={\bm{u}}_{k}+\Delta tf_{\theta}({\bm{u}}_{k}),$ where $f_{\theta}$ is a parametric model of the continuous time dynamics. We parameterize $f_{\theta}$ by a multi-layer perceptron with two hidden layers each of dimension 32 and use the ReLU activation function. We trained with an ADAM optimizer with learning rate $1\mathrm{e}^{-4}.$
Models were trained for 500,000 steps. For the curriculum training (and its regularized counterpart), we increase $\ell$ by one every 50,000 training steps, and hence by the end of training ${\mathcal{S}}_{\theta}$ is predicting trajectories of length 10. For curriculum training, we weight rollout loss using a geometric weighting $\omega(k)=\max(0.1^{k-1},1\mathrm{e}^{-7}).$ For pushforward training we use the same rollout schedule as in curriculum training, but the rollout loss weight is $\omega(k)=\max(0.1^{k-1},1\mathrm{e}^{-4}).$ These weighting schemes were chosen empirically to ensure that training loss was of the same order of magnitude throughout training, even as rollout length increased. The MMD bandwidth values used were $\mathbf{\sigma}=[0.2,0.5,0.9,1.3]$.
@@ E.2 Kuramoto–Sivashinsky
The KS system is known to be chaotic (Papageorgiou & Smyrlis, 1991) and, when stochastically forced, ergodic with an invariant measure (Ferrario, 2008; Weinan & Liu, 2002). We generate data for this system using a spectral solver (Dresdner et al., 2022) on a spatial grid $[0,64]$ with $512$ equally-spaced points and a time a 4th-order implicit-explicit Crack-Nicolson Runge-Kutta scheme (Canuto et al., 2007), with a time resolution of $\Delta t=0.001$ time units. For each trajectory we start with a randomly generated initial condition given by
@@
where $\omega_{j}$ is chosen randomly from $\{2\pi/L,4\pi/L,6\pi/L\}$, $a_{j}$ is sampled from a uniform distribution in $[-0.5,0.5]$, and phase $\phi_{j}$ follows a uniform distribution in $[0,2\pi]$. We use $n_{c}=30$. We let the system to “warm up” for $20$ units of time, before recording the trajectories. The training dataset consists of 800 trajectories of 1,200 steps with a time sampling rate $\Delta t=0.2$ time units, from which we randomly sample batches of size $128$ and trajectory length of $10$ steps. Our evaluation set consists of 100 trajectories of length 1,000 steps.
We parameterize ${\mathcal{S}}_{\theta}$ as a dilated convolution neural network with residual connections, as described in (Stachenfeld et al., 2022). In contrast to the Lorenz 63 model, we do not involve the time step $\Delta t$ directly in the computation of the update, instead we use $\tilde{{\bm{u}}}_{k+1}={\mathcal{S}}({\bm{u}}_{k})$. The architecture consists of an encoder convolutional module, four dilated convolutional blocks, and a decoder convolutional module. There exists a residual connection from the encoder to the output of the first dilated convolution block and from the input to the decoder output. The intermediate representations have $48$ channels. The encoder, decoder, and intermediate dilated convolutions use kernels of width 5. Within each dilated convolution block, there are four dilated convolutional layers followed by ReLU non-linear activations, with dilation factor increasing by a multiple of 2 for each layer. Each block has a residual connection to the previous one. The model has a total of $324,433$ parameters. The model was trained using the ADAM optimizer and an initial learning rate of $5\mathrm{e}^{-4}$. A staircase exponential decay learning rate scheduler was used with a decay factor of $0.5$ and decay transitions every $60,000$ steps.
Models were trained for $300,000$ steps with the rollout increased every 60,000 steps, and hence by the end of training ${\mathcal{S}}_{\theta}$ is predicting trajectories of length 5. For both the Curr and Pfwd objectives we use the same rollout schedule and rollout loss weight: $\omega(k)=\max(0.9^{k-1},1\mathrm{e}^{-3})$.
@@ E.3 Kolmogorov Flow
where $k_{0}=4$. The forcing only acts in the $y$ coordinate. Following (Kochkov et al., 2021), we add a small drag term to dissipate energy. An equivalent problems is given by its vorticity formulation
@@
where $\omega:=\partial_{x}{\bm{u}}_{y}-\partial_{y}{\bm{u}}_{x}$, which we use for spectral method which avoids the need to separately enforce the incompressibility condition $\nabla\cdot{\bm{v}}=0$. The initial conditions are the same as the ones proposed in (Kochkov et al., 2021).
@@
Equations and 33 and 38 were discretized using a pseudo-spectral discretization, to avoid issues stemming from dispersion errors. Pseudo-spectral methods are known to be dispersion free, due to the exact evaluation of the derivatives in Fourier space, while possessing excellent approximation guarantees (Trefethen, 2000). We use the jax-cfd spectral elements tool box Learning to correct spectral methods for simulating turbulent flows, which leverages the Fast Fourier Transform (FFT) (Cooley & Tukey, 1965) to compute the Fourier transform in space of the field ${\bm{u}}(x,t)$, denoted by $\hat{{\bm{u}}}(t)$, which allows for a very efficient computation of spatial derivatives by diagonal rescaling following $\partial_{x}\hat{{\bm{u}}}_{k}=ik\hat{{\bm{u}}}_{k}$, where $k$ is the wave number. This renders the application and inversion of linear differential operators trivial, since they are simply element-wise operations (Trefethen, 2000).
@@
Equation 38 was discretized with spatial discretization $n_{x}=n_{y}=256$ and a 4th order implicit-explicit Crack-Nicolson Runge-Kutta scheme (Canuto et al., 2007), where we treat the linear part implicitly and the nonlinear one explicitly with $\Delta t=0.001$ using jax-cfd (Dresdner et al., 2022).
For each trajectory we let the solver “warm up” for $50$ units of time, in order for the trajectory to reach the attractor. We further evolve the equation for $120$ units of time, and we sample the trajectories at a rate of $\Delta t=0.1$. Finally, we downsample the trajectories by a factor $4$ in each direction. We repeated the process $128$ times to obtain the training data, and $32$ times for both the evaluation and test data.
@@
For the machine learning experiments we set the learning rate to be $5\cdot 10^{-4}$, and we vary the batch size from $32$ to $512$ (depending of the experiment) in increments of power to two. We use an exponential schedules which halves the learning rate every $72.000$ iterations. We trained the models for up to $720.000$ iterations. Unless otherwise stated the MMD bandwidth used was $\mathbf{\sigma}=[2,5,9,13,20,50,90,120]$. This value was found after a quick hyperparameter tuning on a small dataset.
We parametrize ${\mathcal{S}}_{\theta}$ using a two-dimensional dilated convolutional neural network with residual connections and periodic boundary conditions following (Stachenfeld et al., 2022) the total number of parameters is $6,458,785$. In this case we follow the same unrolling scheme than the KS system, i.e., $\tilde{{\bm{u}}}_{k+1}={\mathcal{S}}({\bm{u}}_{k})$.
@@
For Pfwd training, we consider a rollout schedule that follows the learning rate schedule: every $72.000$ iterations we increase the number of unrolling steps, $\ell$, by one, where $\ell$ increases from $1$ to $10$. The effective number of unrolling step at each training step is sampled uniformly from $1$ to $\ell$.
@@
Given the higher memory requirement, we had to decrease the number of unrolling steps from $10$ to just $5$. Also, depending on the batch size we had to further decrease the number of unrolling steps. In particular for large batch sizes we can not afford unrolling more than 2 time steps. All the other parameters were kept constant.
One-step
For this objective, we use the same set up as experiments above. We halved the learning rate every $72.000$ iterations and the models were trained for $720.000$ iterations while keeping $\ell$ constant and equal to one.
@@
Figure 5: Values of the MMD metric for the baselines and the DySLIM regularization, including a finite-sample estimate of the invariant measure.
Figure 6: Wasserstein-1 distance between the different components of the forcing in Equation 32 at different unrolling times. From top to bottom, $x$, $y$, $z$, and the crossed products $xy$ and $xz$.

### v2

@@ 7 Conclusion
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
@@ Appendix A Relation to the Pushforward trick
We note that the pushforward trick (Brandstetter et al., 2022) can be reformulated using our framework as a weak measure fitting loss. A finite-sample approximation of $\mathbb{E}_{{\bm{u}}_{0}\sim\mu^{*}}[||{\mathcal{S}}_{\theta}(\mathtt{sg}({\mathcal{S}}^{k-1}_{\theta}({\bm{u}}_{0})))-{\bm{u}}_{k}||^{2}]$ is an upper bound of the discrete Wasserstein-2 distance between $\mu^{\star}$ and the approximation of $\mu_{\theta}^{\star}$. Formally, we have that
@@
By relying on an estimate of the loss, we have that for a given set of initial conditions $\{{\bm{u}}^{(i)}\}_{i=1}^{n}\sim\mu^{*}$,
@@
which can be lower bounded by the following
@@
where $\pi$ is a permutation operator. Given that we are in the discrete setting where the Monge and Kantorovich problems are equivalent (Brezis, 2018), we have that
@@
where $\Pi$ is the set of all valid discrete transport maps (i.e., matrices that satisfy $T_{i,j}\geq 0$, $\sum_{j}T_{i,j}=\sum_{i}T_{i,j}=1$), $C$ is the quadratic cost function ($C_{i,j}=\|{\mathcal{S}}^{k}({\bm{u}}^{(i)})-{\mathcal{S}}_{\theta}^{k}({\bm{u}}^{(j)})\|^{2}$), and $\widehat{W}_{2}$ is a discrete estimate of the Wasserstein-2 metric.
@@ Appendix B Maximum Mean Discrepancy
In this section, we provide additional information and context about the Maximum Mean Discrepancy (MMD). The MMD is an instance of an integral probability metric (IPM; Müller (1997), which is a useful construction that allows us to measure distance between distributions. For any two distributions $\mu$ and $\nu,$ IPMs are defined with a function class ${\mathcal{G}}$ as:
@@
As described in Section 5, we use a rational quadratic kernel. While other works that use MMD for distribution matching, such as Li et al. (2015) and Dziugaite et al. (2015), also explored the squared exponential kernel, $\kappa_{\sigma}({\bm{u}},{\bm{v}})=\exp(\frac{-1}{2\sigma}||{\bm{u}}-{\bm{v}}||_{2}^{2})$, they found that careful tuning of the bandwidth parameter was required. In contrast, other than the highest dimension Kolmogorov flow experiments, we found that the mixture of bandwidths used in our rational kernel was comparatively robust and did not require a comprehensive hyperparameter search. We therefore rely on this kernel and do not explore the more sensitive squared exponential kernel.
@@ C.1 Cosine Similarity
Letting $\{{\bm{u}}^{(i)}_{t_{k}}\}_{i=1}^{n}$ and $\{\tilde{{\bm{u}}}^{(i)}_{t_{k}}\}_{i=1}^{n}$ be the ground truth and predicted states (respectively) at time $t_{k}$, for $k=1,...,N$, across test set trajectories, the cosine similarity at each time step is defined as:
@@ C.2 Sinkhorn Divergence
Popular metrics used to measure distance between distributions include Optimal Transport (OT) based metrics, such as the Sinkhorn divergence, which we describe below. The field of OT is concerned with transforming (or transporting) one distribution into another, i.e., finding a map between them, in an optimal manner with respect to a pre-defined cost. The cost of the minimal (or optimal) transformation, often called the cost of the OT map, can then be used to define distances between distributions that ‘lifts’ the underlying metric $\mathrm{d}$ defined on ${\mathcal{U}}$ to one over the space of probability measures ${\mathcal{P}}({\mathcal{U}})$ (Santambrogio, 2015).
@@
In Section 5, the SD was used to compare empirical version of the ground truth and predicted distributions of trajectories. We use the Optimal Transport Tools library (Cuturi et al., 2022) with its default hyperparameters to perform this computation. We also explored using the Sinkhorn Divergence as the measure distance in Equation 13. However, especially in higher dimension experiments, we found this divergence to be less informative in guiding training, likely owing to its less favorable estimation properties compared to the MMD, particularly in the high-dimensional regime, see Appendix G for more details.
@@ C.3 Radially Averaged Energy Spectrum
The energy spectrum is one of the main metrics to quantitatively assess generated samples (Wan et al., 2023a). In a nutshell, the energy spectrum measures the energy in each Fourier mode, thereby providing insights into the similarity between the generated and reference samples.
The energy spectrum is defined [^9] This definition is applied to each sample and averaged to obtain the metric (same for MELR). as
@@
where ${\bm{u}}$ is a snapshot system state, $K$ is the magnitude of the wave-number (wave-vector in 2D) $\underline{K}$, and $x_{i,j}$ is the underlying (possibly 2D) spatial grid. To assess the overall consistency of the spectrum between the generated and reference samples using a single scalar measure, we consider the mean energy log ratio (MELR):
@@ C.4 Covariance RMSE (covRMSE)
$\text{covRMSE}=\frac{\|\text{Cov}_{\text{pred}}-\text{Cov}_{\text{ref}}\|}{\|\text{Cov}_{\text{ref}}\|},$ (25)
@@ C.5 Time Correlation Metric (TCM)
can be interpreted as the time for the signal to forget its past. We compute the average pixel-wise $\tau$ for ground truth as well as prediction rollouts and take their absolute difference to form a metric.
@@ Appendix D Regularization
For better reproducibility of our work, we provide explicit formulas for the regularized objective functions. We reproduce Equation 13 from Section 5
@@ Appendix E Experimental Setup
Below, we provide information about each dynamical system from Section 5 and their corresponding experimental setup. In Table 4, we give an overview of the model, learning rate, and number of training steps used in each experiment.
Table 4: Model, learning rate, and number of training steps for each experiment in Section 5.
System ${\mathcal{S}}_{\theta}$ LR Training steps
Lorenz 63 MLP w/residual connection to input $1\mathrm{e}^{-4}$ 500k
Kuramoto–Sivashinsky Dilated convolutional network (Stachenfeld et al., 2022) $5\mathrm{e}^{-4}$ 300k
Kolmogorov Flow Dilated convolutional network (Stachenfeld et al., 2022) $5\mathrm{e}^{-4}$ 720k
@@ E.1 Lorenz 63
The Lorenz 63 model (Lorenz, 1963) is defined on a 3-dimensional state space by the following non-linear ordinary differential equation $\dot{{\bm{u}}}=f({\bm{u}})$:
@@
The Lorenz 63 system is typically associated to parameter values of $\sigma=10,\rho=28,$ and $\beta=8/3$ and is known to be chaotic with an attractor that supports an ergodic measure (Tucker, 2002; Luzzatto et al., 2005).
@@
We define ${\mathcal{S}}_{\theta}$ as a one-step finite difference model: $\tilde{{\bm{u}}}_{k+1}={\mathcal{S}}_{\theta}({\bm{u}}_{k})={\bm{u}}_{k}+\Delta tf_{\theta}({\bm{u}}_{k}),$ where $f_{\theta}$ is a parametric model of the continuous time dynamics. We parameterize $f_{\theta}$ by a multi-layer perceptron (MLP) with two hidden layers each of dimension 32 and use the ReLU activation function. We trained with an ADAM optimizer (Kingma & Ba, 2014) with learning rate $1\mathrm{e}^{-4}.$
Models were trained for 500,000 steps. For the curriculum training (and its regularized counterpart), we increase $\ell$ by one every 50,000 training steps, and hence by the end of training ${\mathcal{S}}_{\theta}$ is predicting trajectories of length 10. For curriculum training, we weight rollout loss using a geometric weighting $\omega(k)=\max(0.1^{k-1},1\mathrm{e}^{-7}).$ For pushforward training we use the same rollout schedule as in curriculum training, but the rollout loss weight is $\omega(k)=\max(0.1^{k-1},1\mathrm{e}^{-4}).$ These weighting schemes were chosen empirically to ensure that training loss was of the same order of magnitude throughout training, even as rollout length increased. The MMD bandwidth values used were $\bm{\sigma}=\{0.2,0.5,0.9,1.3\}$.
@@ E.2 Kuramoto–Sivashinsky
The KS system is known to be chaotic (Papageorgiou & Smyrlis, 1991) and, when stochastically forced, ergodic with an invariant measure (Weinan & Liu, 2002; Ferrario, 2008). We generate data for this system using a spectral solver (Dresdner et al., 2022) on a spatial grid $[0,64]$ with $512$ equally-spaced points and a 4th-order implicit-explicit Crack-Nicolson Runge-Kutta scheme (Canuto et al., 2007), with a time resolution of $\Delta t=0.001$. For each trajectory, we start with a randomly generated initial condition given by
@@
where $\omega_{j}$ is chosen randomly from $\{2\pi/L,4\pi/L,6\pi/L\}$, $a_{j}$ is sampled from a uniform distribution on $[-0.5,0.5]$, and phase $\phi_{j}$ follows a uniform distribution on $[0,2\pi]$. We use $n_{c}=30$. We let the system “warm up” for 20 units of time, before recording the trajectories. The training dataset consists of 800 trajectories of 1,200 steps with a time sampling rate $\Delta t=0.2$ time units, from which we randomly sample batches of size 128 and trajectory length of 10 steps. Our evaluation set consists of 100 trajectories of length 1,000 steps.
We parameterize ${\mathcal{S}}_{\theta}$ as a dilated convolution neural network with residual connections, as described in (Stachenfeld et al., 2022). In contrast to the Lorenz 63 model, we do not involve the time step $\Delta t$ directly in the computation of the update, instead we use $\tilde{{\bm{u}}}_{k+1}={\mathcal{S}}_{\theta}({\bm{u}}_{k})$. The architecture consists of an encoder convolutional module, four dilated convolutional blocks, and a decoder convolutional module. There exists a residual connection from the encoder to the output of the first dilated convolution block and from the input to the decoder output. The intermediate representations have $48$ channels. The encoder, decoder, and intermediate dilated convolutions use kernels of width 5. Within each dilated convolution block, there are four dilated convolutional layers followed by ReLU non-linear activations, with dilation factor increasing by a multiple of 2 for each layer. Each block has a residual connection to the previous one. The model has a total of 324,433 parameters. The model was trained using the ADAM optimizer and an initial learning rate of $5\mathrm{e}^{-4}$. A staircase exponential decay learning rate scheduler was used with a decay factor of 0.5 and decay transitions every 60,000 steps.
Models were trained for 300,000 steps with the rollout increased every 60,000 steps, and hence by the end of training ${\mathcal{S}}_{\theta}$ is predicting trajectories of length 5. For both the Curr and Pfwd objectives we use the same rollout schedule and rollout loss weight: $\omega(k)=\max(0.9^{k-1},1\mathrm{e}^{-3})$. The MMD bandwidth values used were $\bm{\sigma}=\{0.2,0.5,0.9,1.3\}$.
@@ E.3 Kolmogorov Flow
where $k_{0}=4$. The forcing only acts in the $y$ coordinate. Following Kochkov et al. (2021), we add a small drag term to dissipate energy. An equivalent problem is given by its vorticity formulation
@@
where $\omega:=\partial_{x}{\bm{u}}_{y}-\partial_{y}{\bm{u}}_{x}$, which we use for spectral method which avoids the need to separately enforce the incompressibility condition $\nabla\cdot{\bm{v}}=0$. The initial conditions are the same as the ones proposed in Kochkov et al. (2021).
@@
Equations 33 and 38 were discretized using a pseudo-spectral discretization, to avoid issues stemming from dispersion errors. Pseudo-spectral methods are known to be dispersion free, due to the exact evaluation of the derivatives in Fourier space, while possessing excellent approximation guarantees (Trefethen, 2000). We use the jax-cfd spectral elements tool box (Dresdner et al., 2022). Learning to correct spectral methods for simulating turbulent flows, which leverages the Fast Fourier Transform (Cooley & Tukey, 1965) to compute the Fourier transform in space of the field ${\bm{u}}(x,t)$, denoted by $\hat{{\bm{u}}}(t)$, allows for a very efficient computation of spatial derivatives by diagonal rescaling following $\partial_{x}\hat{{\bm{u}}}_{K}=iK\hat{{\bm{u}}}_{K}$, where $K$ is the wave number. This renders the application and inversion of linear differential operators trivial, since they are simply element-wise operations (Trefethen, 2000).
@@
Equation 38 was discretized with spatial discretization $n_{x}=n_{y}=256$ and a 4th order implicit-explicit Crack-Nicolson Runge-Kutta scheme (Canuto et al., 2007), where we treat the linear part implicitly and the nonlinear one explicitly with $\Delta t=0.001$ using jax-cfd.
For each trajectory, we let the solver “warm up” for 50 units of time, in order for the trajectory to reach the attractor. We further evolve the equation for 120 units of time, and we sample the trajectories at a rate of $\Delta t=0.1$. Finally, we downsample the trajectories by a factor 4 in each spatial direction. We repeated the process 128 times to obtain the training data, and 32 times for both the validation and test data.
@@
We set the learning rate to be $5\mathrm{e}^{-4}$, and we vary the batch size from 32 to 512 (depending of the experiment) in increments of power to two. We use an exponential learning rate scheduler, which halves the learning rate every 72,000 iterations. We trained the models for up to 720,000 iterations. Unless otherwise stated the MMD bandwidth used was $\bm{\sigma}=\{2,5,9,13,20,50,90,120\}$. This value was found after a quick hyperparameter tuning on a small dataset.
We parametrize ${\mathcal{S}}_{\theta}$ using a two-dimensional dilated convolutional neural network with residual connections and periodic boundary conditions following (Stachenfeld et al., 2022) the total number of parameters is 6,458,785. We follow the same unrolling scheme as in the KS system, i.e., $\tilde{{\bm{u}}}_{k+1}={\mathcal{S}}({\bm{u}}_{k})$.
One-step
For this objective, we use the same set up as experiments above. We halved the learning rate every 72,000 iterations and the models were trained for 720,000 iterations while keeping $\ell$ constant and equal to one.
@@
For Pfwd training, we consider a rollout schedule that follows the learning rate schedule: every 72,000 iterations we increase the number of unrolling steps $\ell$ by one, where $\ell$ increases from $1$ to $10$. The effective number of unrolling step at each training step is sampled uniformly from $1$ to $\ell$.
@@
Given the higher memory requirement, we decrease the number of maximum unrolling steps from $10$ to just $5$. Also, depending on the batch size, we further decrease the maximum number of unrolling steps. In particular, for large batch sizes, we cannot afford unrolling more than 2 time steps. All the other parameters were kept constant relative to Pushfoward experiments.
@@
Figure 5: Values of the MMD metric for the baselines and the DySLIM regularization. Each point represents a random training seed that remains stable, with the solid line indicating median values.
Figure 6: Wasserstein-1 distance ($\downarrow$) between the different components of the forcing in Equation 32 at different unrolling times. From top to bottom, $x$, $y$, $z$, and the crossed products $xy$ and $xz$. Each point represents a random training seed that remains stable, with the solid line indicating median values.